Find the equation of a line described as follows and express your answer in point-slope form, slope-intercept form, and standard form
Find the equation of the line perpendicular to the line 2x-3y=5 that passes through the point (1, -4)

Answer by ewatrrr(24785)

y = (2/3)x - 5/3
Line perpendicular: m = -3/2
point-slope form, y +4 = (-3/2)(x - 1)
slope-intercept form, y = (-3/2)x -5/2
and standard form 3x + 2y = -5